Calculate equivalent conductivity of 1 M H2SO4 whose conductivity is 26 × 10−2 ohm−1 cm−1
    (A)    260
    (B)    130
    (C)    65
    (D)    200

Grade:upto college level

2 Answers

Radhika Batra
10 years ago

Sp. conductance = 26 × 10−2 Ω cm−1

    Concentration = 1 MH2SO4

            = 98 g/L

    Eq. wt. of H2SO4 = 49

    ? No. of gram equivalents per litre = 98/49=2

    Equivalent conductance = (K×1000)/c

                    =(26×?10?^(-2)×1000)/2

                    =130 Ω cm2 mol−1

Rishi Sharma
3 years ago
Dear Student,
Please find below the solution to your problem.

Sp. conductance = 26 × 10−2 Ω cm−1
Concentration = 1 MH2SO4 = 98 g/L
Eq. wt. of H2SO4 = 49
No. of gram equivalents per litre = 98/49=2
Equivalent conductance = (K×1000)/c
=(26×10^(-2)×1000)/2
=130 Ω cm2 mol−1

Thanks and Regards